# appReferences in the Appendix

### v2

@@ 5 Conclusion
Structure of the Appendix
We organize the appendix as follows. In Section A, we introduce the foundation analysis for LP-based method and first-order method for the discussion of our framework; Section B presents the cohesive mathematical framework which unify all the OLP methods, and the unique properties derived from our parallel multi-phase algorithm; Section C combines all the previous results and proves the main results in our paper; Section D provides some auxiliary results and technical support in the previous three parts.
Appendix A Primary Results and Properties
In this section, we provide primary results of LP-based method and first-order method. We also include some basic analysis of dual algorithms that will help present our main results. We start from the convergence of regular LP-based method. All the analysis are under Assumption (1) and (2).
A.1 LP-based Analysis
We provide the analysis for dual solution obtained from LP-based method, so $\bm{p}_{t}=\bm{p}_{t}^{L}$ in this part.
@@ A.1 LP-based Analysis
Proof. By the optimality of $\bm{p}^{*}$ and boundedness of $\bm{d}$, we have
@@
Similarly, by the optimality of $\bm{p}_{t}^{*}$ and its associated $\bm{d}_{t}\in\mathcal{D}$ in Lemma 2.1, we know
@@
Proof of Lemma 3.2.
According to the latest results in [13], the online dual price $\bm{p}_{t}$ achieves a sublinear convergence $\mathcal{O}(\frac{1}{\sqrt{t}})$ to $\bm{p}_{t}^{*}$. Since
@@
By Lemma 12 in [16], we have $\|\bm{p}_{t}^{*}-\bm{p}^{*}\|_{2}^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\|\bm{d}_{t}-\bm{d}\|_{2}^{2}$.
A.2 First-order Analysis
We provide the analysis for dual solution obtained from LP-based method, so $\bm{p}_{t}=\bm{p}_{t}^{D}$ in this part.
@@ A.2 First-order Analysis
Proof. As our initial choice $\bm{p}_{1}=\bm{0}$, according to Lemma 1 in [15], we get the above result. In addition, by Lemma B.1 in [8], we have
@@
2) With quadratic regularity of $f$ as in Lemma 2.2, we have:
@@
Case 2. When $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, for any $j\leq t$, Lemma 3.3 gives us
@@ Appendix B Algorithm Design and Analysis
In this section, we present the cohesive mathematical framework to integrate the LP-based method and first-order method by developing a new performance metric. We also provide some unique properties for our algorithm design of parallel multi-phase structure.
As stated in Algorithm 3, we have independent two paths for online learning and online decision-making conducted simultaneously. Specifically, for the online learning path, we use the LP-based method to resolve the updated linear programs at a fixed frequency and update this result to the decision-making path; For online decision-making path, we restart this process every time when receiving an updated learning result. We apply first-order method in the initial and final batches, and use the latest dual price for decision-making in the intermediate resolving intervals. This process is illustrated in Figure (1).
B.1 Optimality Gap Analysis
@@
Under Assumptions (1) and (2), define the deterministic Lagrangian dual function as
@@
Proof. This lemma is proved as Lemma 3 in [16] by strong duality and optimality of $\bm{p}_{T}^{*}$.
Theorem B.1 (Decomposition of Optimality Gap).
Under Assumptions (1) and (2), let $C_{r}=\max\{\frac{\bar{r}}{\underline{d}},\frac{\nu\lambda\bar{a}^{2}}{2}\}$, we derive an upper bound for the optimality gap $r(\bm{x})$ as:
@@
Proof. By the definition of $\ell(\bm{p})$, we derive the online objective as
@@
Then by Lemma B.1, we derive the upper bound for the optimality gap $r(\bm{x})$ as defined in (8) to be:
@@
B.2 Total Regret Analysis
We consider the constraint violation as $v(\bm{x})=\lVert(\bm{Ax}-\bm{b})^{+}\rVert$ in (9). Then the total regret is the combination of optimality gap and constraint violation. Since Theorem B.1 holds for any unknown distribution $\mathcal{P}\in\Xi$, we have:
@@ B.2 Total Regret Analysis
This new framework is adaptive to all OLP methods. We will proceed based on this structure to analyze various methods respectively.
@@
Based on the change of methods, we separate the horizon into three intervals $T_{1},T_{2}$, and $T_{3}$, and reorganize the total regret for analysis to be
@@
where $\Delta_{T_{1}},\Delta_{T_{2}}$, and $\Delta_{T_{3}}$ are the total regret during each interval.
Proof. We split the total horizon into three intervals of initial batch $T_{1}=[0,f]$, intermediate process $T_{2}=[f,kf]$, and final batch $T_{3}=[kf,T]$. We use the first-order method for the first and final batches and LP-based method for intermediate process. Notice that for the initial resolving batch, we maintain the classical analysis for first-order method since it has not reached to the mixture with LP-based method. Then based on (B.2), we define
@@ B.3 Regret for LP-based Method
In this section, we are going to bound $\Delta_{T_{2}}$ for the regret from LP-based method. To provide a tighter analysis, we examine the real-time average resource capacity $\bm{d}_{t}$ instead of the original process $\bm{b}_{t}$. Specifically, by Lemma 2.1, we constraint $\bm{d}_{t}$ within a feasible range $\mathcal{D}$. If $\bm{d}_{t}\notin\mathcal{D}$, we set the dual price to zero to accept all subsequent orders. We analyze the remaining resource and constraint violation due to this strategy. Note that this is required solely for rigor analysis purpose and is not necessary when running the algorithm.
@@
Under Assumption 1 and 2, there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda_{1},\mu$, and $C_{lp}$ such that
@@
Proof. During the re-solving process, we denote each interval to be $[jf,(j+1)f]$ where $j=1,2,...,k$. As the resource usage follows $\bm{b}_{t+1}=\bm{b}_{t}-\bm{a}_{t+1}I(r_{t+1}>\bm{a}_{t+1}^{T}\bm{p}_{{t+1}})$, normalizing both sides, we derive the update of average resource consumption to be:
@@
Theorem B.3 (Regret of Intermediate Process).
Under Assumption 1 and 2, following the result in (B.2), we prove the regret satisfies
$\Delta_{T_{2}}\leq\log\Big{(}\frac{T}{f}\Big{)}=\mathcal{O}(\log(k)).$ (37)
Proof. We analyze the three terms in (B.2) respectively.
@@
    For dual convergence, by Lemma 3.2, we have
@@ B.4 Regret for First-Order Method
In this section, we are going to bound $\Delta_{T_{1}}$ and $\Delta_{T_{3}}$ for the regret from first-order method. The key point is to analyze the connection between dual prices based on the gradient-updated rule. Here $\bm{p}_{t}=\bm{p}_{t}^{fo}$.
@@
Under Assumption 1 and 2, we have
@@
Proof. We analyze $\Delta_{T_{1}}$ in (B.2) by parts. By Lemma B.3, B.4 in [8], we have
$\sum_{t=1}^{f}\mathbb{E}[r_{t}x_{t}^{*}-r_{t}x_{t}]\leq\frac{m(\bar{a}+\bar{d})^{2}}{2}\alpha_{t}f,$
$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=1}^{f}\bm{a}_{t}x_{t}-f\bm{d}\Big{)}^{+}\Big{\rVert}\leq\frac{1}{\alpha_{t}}\bigg{[}\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\bigg{]}.$
To achieve the tight upper bound, we select the optimal step size $\alpha_{t}=\frac{1}{\sqrt{f}}$, which then gives us the desired result.
@@
Under Assumption 1 and 2, take the learning rate $\alpha_{t}=\alpha$, we bound the constraint violation for the last batch as
@@
Proof. With the update rule of first-order method in Algorithm 2, we have
@@
Summarizing on both sides and apply the telescoping, we derive
@@
Thus, according to Lemma 3.2 and (46), the expectation of constraint violation satisfies
@@
Theorem B.5 (Regret of Final Batch).
Under Assumption 1 and 2, we have
Proof. We decompose $\Delta_{T_{3}}$ into three parts according to (B.2) and analyze each term respectively. Since the first-order method re-starts from $\bm{p}_{kf}$, by Lemma 3.2, we have:
@@
Since $t\in[kf,T]$, by Proposition 3.3 in [8], the dual price is far from the origin and thus the positive projection still equals to itself. This indicates that:
@@
Summarizing on both sides and apply the telescoping, we derive
@@
Then for the positive binding terms, by Lemma 3.2 and (46), we have
@@
    If $T-kf\leq f^{1/3}$, the total regret must be smaller than the length of the batch, so $\Delta_{T_{3}}\leq T-kf\leq f^{1/3}$.
@@
Lemma B.4 (Warm-Start First-Order Regret).
Assume that we have a warm start for the initial dual price $\bm{p}_{0}$ in the first batch, which satisfies $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$. Then the regret of first batch follows
Proof. According to Lemma D.3, with $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ and $\alpha<\frac{1}{\nu\lambda}$, we have
@@
Thus, take the optimal learning rate $\alpha=f^{-2/3}$, we obtain the regret for the first batch with warm start as:

### v3

@@ A.1 Preliminary results
This lemma is a consequence of Lemmas 12 and 13 in (Li & Ye, 2022). The existence of $\delta$ comes from the continuity of $f_{\bm{d}_{t}}(\bm{p})$. Thus, $\delta$ is only associated with stochastic program (7), and it is independent of $T$. Note that $\mathcal{D}\subset\mathcal{V}_{\bm{d}}$. We will analyze our feasible dual solutions derived from the online algorithm with $\bm{d}_{t}\in\mathcal{D}$.
Lemma A.2 (Quadratic Regularity, Proposition 2 in (Li & Ye, 2022)).
Under Assumptions 2.1 and 2.2, for any $\bm{p}\in\mathcal{V}_{\bm{p}}$,
Moreover, $\bm{p}^{*}$ is the unique optimal solution to (2).
A.2 LP-based analysis
We state the LP-based method in Algorithm˜3 and its properties of boundedness and dual convergence results.
Algorithm 3 LP-based method
for $t$ = $1$ to $T$ do
   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6) if constraints are not violated. Update constraint for i=1,,mi=1,...,m:   remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$   average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Solve the updated dual problem and obtain dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
end for
LP-based method incorporates the past decisions into the optimization of $\bm{p}_{t}$. If resources were over-utilized in earlier periods, the remaining supply decreases, prompting the algorithm to raise the dual price and become more selective with future orders. Conversely, if ample resources remain, the future dual price will be lowered, allowing more consumer requests to be accepted. This adaptive mechanism accounts for past actions by adjusting the available resource capacity.
@@ A.2 LP-based analysis
Lemma A.4 (Dual Convergence of LP-based algorithm).
Under Assumptions 2.1 and 2.2, $\bm{p}_{t}$ represents the online solution from LP-based method, there exists a constant $C_{lp}>0$ depending on $\bar{r},\bar{a},\underline{d},m,\nu$, and $\lambda$ such that
$\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}_{t}^{*}\rVert^{2}\right]\leq\frac{C_{lp}}{t}.$ (15)
In addition, the difference between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ satisfies
This lemma establishes that the LP-based online AHDL algorithm (Jiang et al., 2022) produces dual solutions $\bm{p}_{t}$ that converge to $\bm{p}_{t}^{*}$. This convergence highlights the stability of the online dual variables during the intermediate stages of decision-making and ensures a warm start for the first-order method in the last re-solving batch. Moreover, we bound the distance between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ by relating their respective average resource capacities, $\bm{d}_{t}$ and $\bm{d}$, in the associated stochastic programs (Li & Ye, 2022). These results provide the foundation for analyzing $\lVert\bm{p}_{t}-\bm{p}^{*}\rVert$ in our final result.
Proof of ˜A.3.
By the optimality of $\bm{p}^{*}$ and boundedness of $\bm{d}$, we have
@@
Similarly, by the optimality of $\bm{p}_{t}^{*}$ and its associated $\bm{d}_{t}\in\mathcal{D}$ in Lemma A.1, we know
@@
Proof of ˜A.4.
According to the latest results in (Jiang et al., 2022), the online dual price $\bm{p}_{t}$ achieves a sublinear convergence $\mathcal{O}(\frac{1}{\sqrt{t}})$ to $\bm{p}_{t}^{*}$. Since
@@
By Lemma 12 in (Li & Ye, 2022), we have $\|\bm{p}_{t}^{*}-\bm{p}^{*}\|_{2}^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\|\bm{d}_{t}-\bm{d}\|_{2}^{2}$.
A.3 First-order Analysis
We present the classic first-order method (Li et al., 2020) in Algorithm˜4 and the first-order method with restart strategy (Gao et al., 2024) in Algorithm˜5. We also show their properties of boundedness and dual convergence results.
Algorithm 4 First-Order Online algorithm
for $t$ = $1$ to $T$ do
   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (5). Update learning rate $\alpha_{t}=\mathcal{O}(1/\sqrt{T})$. Compute subgradient and obtain dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ (\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}+\frac{1}{2\alpha_{t}}\lVert\bm{p}-\bm{p}_{t}\rVert^{2}$
end for
Algorithm 5 First-Order Restart algorithm
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0},\bm{p}^{L}_{1}=\bm{0}$
for $t$ = $1$ to $T$ do
   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6) Update learning rate $\alpha_{t}=\begin{cases}\mathcal{O}(1/T^{1/3})&t\leq T^{2/3}\\ \mathcal{O}(1/T^{2/3})&t>T^{2/3}\end{cases}$ Compute subgradient and obtain dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ (\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}+\frac{1}{2\alpha_{t}}\lVert\bm{p}-\bm{p}_{t}\rVert^{2}$
Run subgradient method with stepsize $\alpha_{t}=\mathcal{O}(1/t)$ and update $\{\bm{p}_{t}^{L}\}$. At $t=T^{2/3}$, restart $\bm{p}_{T^{2/3}}=\bm{p}^{L}_{T^{2/3}}$.
end for
@@ A.3 First-order Analysis
Lemma A.6 (Dual convergence of First-order algorithm).
Under Assumptions 2.1 and 2.2, $\bm{p}_{t}$ represents the online solution from first-order method, if $\alpha_{t}<\nu\lambda$, the subgradient updates satisfy the following recursion rule:
$\mathbb{E}\left[\lVert\bm{p}_{t+1}-\bm{p}^{*}\rVert^{2}\right]$ $\leq(1-\alpha_{t}\nu\lambda)\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}.$ (17)
$\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]\leq C_{fo}\left(\frac{1}{\alpha t}+\alpha\right).$ (18)
$\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]\leq\frac{C_{fo}}{t}.$ (19)
This lemma shows that the first-order method (Gao et al., 2024) guarantees the convergence of $\bm{p}_{t}$ to $\bm{p}^{*}$. It also highlights a key trade-off in choosing the learning rate $\alpha_{t}$, an aspect that will be central to our later optimality analysis. It also highlights a key trade-off in choosing the learning rate $\alpha_{t}$, an aspect that will be central to our later optimality analysis. With this groundwork in place, we now move on to bounding the total regret of our algorithm.
Proof of ˜A.5.
As our initial choice $\bm{p}_{1}=\bm{0}$, according to Lemma 1 in (Li et al., 2020), we get the above result. In addition, by Lemma B.1 in (Gao et al., 2024), we have
@@
2) With quadratic regularity of $f$ as in Lemma A.2, we have:
@@
Case 2. When $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, for any $j\leq t$, Lemma A.6 gives us
@@ Appendix B Algorithm Design and Analysis
In this section, we present a cohesive mathematical framework to integrate the LP-based method and first-order method by developing a new performance metric. We also provide some unique properties for our algorithm design of parallel multi-phase structure.
As stated in Algorithm 1, we construct independent paths for online learning and online decision-making conducted simultaneously. Specifically, for the online learning path, we use the LP-based method to resolve updated linear programs at a fixed frequency and send this result to the decision-making path; For the online decision-making path, we apply the first-order method in the initial and final batches, and use the latest dual price for decision-making in the intermediate resolving intervals. This process is illustrated in Figure˜1.
B.1 Regret Analysis
@@
Under Assumptions (2.1) and (2.2), define the deterministic Lagrangian dual function as
@@
Lemma B.2 (Dual Price Boundedness).
Under Assumption 2.1, the online and offline optimal dual prices are bounded respectively by $\lVert\bm{p}^{*}\rVert\leq\tfrac{\bar{r}}{\underline{d}}$ and
$\lVert\bm{p}_{t}\rVert\leq\bar{p}\coloneqq\max\bigg(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\bigg).$
This lemma establishes that the optimal dual prices remain bounded. Our algorithm maintains these bounds because if $\bm{p}_{t}$ grows large, the algorithm responds by accepting more orders, which in turn reduces $\bm{p}_{t+1}$. This self-correcting mechanism keeps the dual prices within these limits. With this groundwork in place, we now move on to derive the upper bound for the total performance of Algorithm˜1.
Theorem B.3 (Decomposition of Regret).
Under Assumptions (2.1) and (2.2), let $C_{r}=\max\{\frac{\bar{r}}{\underline{d}},\frac{\nu\lambda\bar{a}^{2}}{2}\}$, we derive an upper bound for the regret $r(\bm{x})$ as:
@@
Proof of ˜B.1.
This lemma is proved as Lemma 3 in (Li & Ye, 2022) by strong duality and optimality of $\bm{p}_{T}^{*}$.
Proof of Lemma B.2.
By the boundedness results in Lemma A.3 and Lemma A.5, define
then we complete the proof.
Proof of ˜B.3.
By the definition of $\ell(\bm{p})$, we derive the online objective as
@@
Then by Lemma B.1, we derive the upper bound for the regret $r(\bm{x})$ as defined in (8) to be:
@@
B.2 Total Performance Analysis
We consider the constraint violation as $v(\bm{x})=\lVert(\bm{Ax}-\bm{b})^{+}\rVert$ in (9). Then the total performance is the combination of regret and constraint violation. Since Theorem B.3 holds for any unknown distribution $\mathcal{P}\in\Xi$, we have:
@@ B.2 Total Performance Analysis
This new framework is adaptive to all OLP methods. We proceed with this structure to analyze various OLP methods.
@@
Based on the change of methods, we separate the horizon into three intervals $T_{1},T_{2}$, and $T_{3}$, and reorganize the total performance for analysis to be
@@
where $\Delta_{T_{1}},\Delta_{T_{2}}$, and $\Delta_{T_{3}}$ are the performances for each interval.
Proof of ˜B.4.
We split the total horizon into three intervals of initial batch $T_{1}=[0,f]$, intermediate process $T_{2}=[f,kf]$, and final batch $T_{3}=[kf,T]$. We use the first-order method for the first and final batches and the LP-based method for intermediate processes. Notice that for the initial resolving batch, we maintain the classical analysis for the first-order method since it has not reached the mixture with the LP-based method. Then based on (B.2), we define
@@ B.3 Regret for LP-based Method
In this section, we are going to bound $\Delta_{T_{2}}$ for the regret from LP-based method. To provide a tighter analysis, we examine the real-time average resource capacity $\bm{d}_{t}$ instead of the original process $\bm{b}_{t}$. Specifically, by Lemma A.1, we constraint $\bm{d}_{t}$ within a feasible range $\mathcal{D}$. If $\bm{d}_{t}\notin\mathcal{D}$, we set the dual price to zero to accept all subsequent orders. We analyze the remaining resources and constraint violation due to this strategy. Note that this is required solely for rigor analysis purposes and is not necessary when running the algorithm.
@@
Under Assumption 2.1 and 2.2, there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda_{1},\mu$, and $C_{lp}$ such that
@@
Theorem B.6 (Regret of Intermediate Process).
Under Assumption 2.1 and 2.2, following the result in (B.2), we prove the regret satisfies
$\Delta_{T_{2}}\leq\log\Big(\frac{T}{f}\Big)=\mathcal{O}(\log(k)).$ (26)
Proof of ˜B.5.
During the re-solving process, we denote each interval to be $[jf,(j+1)f]$ where j=1,2,,kj=1,2,...,k. As the resource usage follows $\bm{b}_{t+1}=\bm{b}_{t}-\bm{a}_{t+1}I(r_{t+1}>\bm{a}_{t+1}^{T}\bm{p}_{{t+1}})$, normalizing both sides, we derive the update of average resource consumption to be:
@@
Proof of ˜B.6.
We analyze the three terms in (B.2) respectively.
@@
    For dual convergence, by Lemma A.4, we have
@@ B.4 Regret for First-Order Method
In this section, we are going to bound $\Delta_{T_{1}}$ and $\Delta_{T_{3}}$ for the regret from the first-order method. The key point is to analyze the connection between dual prices based on the gradient-updated rule. Here $\bm{p}_{t}=\bm{p}_{t}^{fo}$.
@@
Under Assumption 2.1 and 2.2, we have
@@
Under Assumption 2.1 and 2.2, take the learning rate $\alpha_{t}=\alpha$, we bound the constraint violation for the last batch as
@@
Theorem B.9 (Regret of Final Batch).
Under Assumption 2.1 and 2.2, we have
Lemma B.10 (Warm-Start First-Order Regret).
Assume that we have a warm start for the initial dual price $\bm{p}_{0}$ in the first batch, which satisfies $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$. Then the regret of the first batch follows
Proof of ˜B.7.
We analyze $\Delta_{T_{1}}$ in (B.2) by parts. By Lemma B.3, B.4 in (Gao et al., 2024), we have
$\sum_{t=1}^{f}\mathbb{E}[r_{t}x_{t}^{*}-r_{t}x_{t}]\leq\frac{m(\bar{a}+\bar{d})^{2}}{2}\alpha f,$
$\mathbb{E}\Big\lVert\Big(\sum_{t=1}^{f}\bm{a}_{t}x_{t}-f\bm{d}\Big)^{+}\Big\rVert\leq\frac{1}{\alpha}\bigg[\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\bigg].$
To achieve the tight upper bound, we select the optimal step size $\alpha=\frac{1}{\sqrt{f}}$, which then gives us the desired result.
Proof of ˜B.8.
With the update rule of first-order method in Algorithm 4, we have
@@
Summarizing on both sides and applying the telescoping, we derive
@@
Thus, according to Lemma A.4 and (41), the expectation of constraint violation satisfies
@@
Proof of ˜B.9.
We decompose $\Delta_{T_{3}}$ into three parts according to (B.2) and analyze each term respectively. Since the first-order method re-starts from $\bm{p}_{kf}$, by Lemma A.4, we have:
@@
Since $t\in[kf,T]$, by Proposition 3.3 in (Gao et al., 2024), the dual price is far from the origin and thus the positive projection still equals itself. This indicates that:
@@
Summarizing on both sides and applying the telescoping, we derive
@@
Then for the positive binding terms, by Lemma A.4 and (41), we have
@@
    If $T-kf\leq f^{1/3}$, the regret must be smaller than the length of the batch, so $\Delta_{T_{3}}\leq T-kf\leq f^{1/3}$.
@@
Proof of ˜B.10.
According to Lemma D.3, with $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ and $\alpha<\frac{1}{\nu\lambda}$, we have
@@
Thus, taking the optimal learning rate $\alpha=f^{-2/3}$, we obtain the regret for the first batch with a warm start as: